# Robust Distributed Gradient Descent to Corruption over Noisy Channels

### v1

@@ V Corruption Reduction by Exploiting the Best of Both Worlds
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{\alpha}{2}\Big{(}1-\frac{\alpha}{M}\Big{)}^{t_{0}-1}\frac{t_{0}(t_{0}+1)}{t(t+1)}\|\theta^{*}-\theta_{0}\|_{2}^{2}\\ +R_{\Theta}t^{r}\cdot A(t,t_{0})+\frac{M^{2}}{2\alpha}+(M+1)\sqrt{p}\sigma R_{\Theta},$
@@ VI-A Least Squares Regression on Synthetic Datasets
The performance metric is the suboptimality gap $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$. The results reported are averaged over 100 independent trials. We also report the standard deviations across these 100 experiments. The bound of the total corruption until time $t$ is set to $C(t)=20t^{0.4}$. As it is difficult to find the worst case corruption pattern, we use the following procedure as a proxy to generate the corruption $\varepsilon_{i,t}$ at each worker $i\in[m]$ and at each time $t$. In particular, let $\Gamma(\tau)=\sum_{k=1}^{\tau}c(k)^{2}$ denote the corruption budget used up to time and including $n$. Then, the available corruption budget at time $t$ is $\sqrt{C(t)^{2}-\Gamma(t-1)}$. Then, we distribute this available corruption budget to each worker randomly. The sign of adversarial corruption at each worker $\operatorname{sign}(\varepsilon_{i,t})$ is set to $-\operatorname{sign}(\tilde{g}_{i,t})$ for all $i\in[m]$. Also, the stepsize is set to $\eta_{t}=t^{-2}$.
@@
As shown in Fig. 4, if one chooses the stepsize sequence $\eta_{k}=\frac{\alpha}{M}H_{k}$, the adversarial corruption accumulates after a certain point, resulting in a stagnation of the suboptimality gap. However, it is also worth noting that the gap decreases much more rapidly when $\eta_{k}=\frac{\alpha}{M}H_{k}$ compared to when $\eta_{k}=\frac{2}{k+1}H_{k}$ initially. In our proposed algorithm with a restart mechanism RDGD-Restart, the gap decreases rapidly until a transition time $t_{0}$ which can be determined analytically. After the time $t_{0}+1$, we switch the stepsize sequence from $\eta_{k}=\frac{\alpha}{M}H_{k}$ to $\eta_{k}=\frac{2}{k+1}H_{k}$. As can be seen, the gap subsequently decreases with $t$. Our proposed algorithm RDGD-Restart converges faster than the $\eta_{k}=\frac{2}{k+1}H_{k}$ case throughout the entire process, indicating a smaller time complexity to achieve the same suboptimality gap. The experimental results presented in Fig. 4 thus corroborate our theoretical analysis in Section V. Hence, our proposed algorithm with a restart mechanism RDGD-Restart  successfully takes advantage of the exponential decrease in the gap until corruption accumulation occurs, then mitigates the adverse effects of the accumulation by switching to another stepsize sequence.
@@ VI-B Classification using L2 Support Vector Machines
As shown in Fig. 5, the proposed RDGD converges rapidly and achieves high accuracy at any level of the corruption budget. However, the accuracy of the classification of the vanilla DGD algorithm is lower than RDGD and is highly sensitive to the level of the corruption budget. In particular, the performance of DGD fails when the corruption budget exceeds a certain level (e.g., $r=0.35$). In Fig. 6, we display the decision boundaries for the first two dimensions of the parameter vectors $\hat{\theta}_{\rm{RDGD}}$ and $\hat{\theta}_{\rm{DGD}}$ obtained by RDGD and DGD respectively when $r=0.35$. Note that $\hat{\theta}_{\rm{RDGD}}$ is almost identical to $\theta^{*}$. However, $\hat{\theta}_{\rm{DGD}}$ deviates significantly from $\theta^{*}$, showing that DGD is far from robust.
@@ VI-C Softmax Classification on the MNIST Dataset
In Fig. 7, we compare RDGD with vanilla DGD under varying levels of corruption budgets $C(T)$, with a fixed noise variance $\sigma^{2}=0.5$ in the communication channels. Specifically, the total corruption bound up to time $T$ is set as $C(T)=100T^{r}$, where $r\in\{0.2,0.3,0.4\}$. The total corruption budget is also uniformly allocated to each iteration and each worker as in the previous subsection. As shown in Fig. 7, RDGD  outperforms the traditional DGD algorithm across all levels of the corruption budget. Furthermore, we observe that RDGD achieves higher accuracy as the value of $r$ and hence, the amount of corruption, decreases. The vanilla DGD algorithm reduces to random guessing (success rate $\approx 10\%$) when the corruption budget exceeds a certain level (i.e., $r=0.4$).
@@ VII Conclusion
We may consider the following as avenues for future research. First, to model the scenario in which some, but not all, workers are compromised, we may redefine $c_{t}$ in (4) as $c_{t}=\sum_{i=1}^{m}\|\varepsilon_{i,t}\|_{2}$. This encourages sparsity among the $\varepsilon_{i,t}$’s for fixed $t$. Second, we may derive information-theoretic lower bounds on the suboptimality gap in terms of the amount of corruption $C(T)$ to ascertain the tightness of the upper bounds.
@@ Appendix B Proof of Lemma 2
$H_{t}L_{t}-H_{t-1}L_{t-1}$ $\geqslant\eta_{t}L(\theta_{t})-\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t}\rangle$
@@
$=\eta_{t}L(\theta_{t})-\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle$
$\;+\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+D_{\psi}(\theta_{t+1},\theta_{t}).$
@@
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})-\eta_{t}L(\theta_{t})\\ +\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle-\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\\ -D_{\psi}(\theta_{t+1},\theta_{t}).$
From RDGD in Algorithm 1 and the convexity of the loss function $L$, we can write $H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})-\eta_{t}L(\theta_{t})$ as:
@@
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t})$
@@ Appendix C Proof of Theorem 1
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t}).$
By the definition of the Bregman divergence and the $\mu$-strong convexity of the function $\psi$, we have $D_{\psi}(\theta_{t+1},\theta_{t})\geqslant\frac{\mu}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$. Hence,
$-\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t})\\ \leqslant\eta_{t}M\|\theta_{t+1}-\theta_{t}\|_{2}-\frac{\mu}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\leqslant\frac{\eta_{t}^{2}M^{2}}{2\mu},$
where the second inequality follows from the fact that $2ab-b^{2}\leqslant a^{2}$.
@@
$\!=\!\bigg{\langle}\frac{1}{m}\Big{(}\!\sum_{i=1}^{m}g^{\prime}_{i,t}\!+\!\sum_{i=1}^{m}\varepsilon_{i,t}\!+\!\sum_{i=1}^{m}w_{k}^{(i)}\Big{)}$
$\hskip 56.9055pt-\frac{1}{m}\!\sum_{i=1}^{m}g_{i,t},\theta^{*}\!-\!\theta_{k+1}\bigg{\rangle}$
$=\frac{1}{m}\bigg{\langle}\!\sum_{i=1}^{m}\!g^{\prime}_{i,t}\!-\!\sum_{i=1}^{m}\!g_{i,t},\theta^{*}\!-\!\theta_{k+1}\!\bigg{\rangle}\!+\!\frac{1}{m}\bigg{\langle}\!\sum_{i=1}^{m}\!\varepsilon_{i,t},\theta^{*}\!-\!\theta_{k+1}\!\bigg{\rangle}$
$\hskip 56.9055pt-\frac{1}{m}\bigg{\langle}\sum_{i=1}^{m}w_{k}^{(i)},\theta^{*}-\theta_{k+1}\bigg{\rangle}.$
@@
$\leqslant MR_{\Theta}\|v_{k}^{(i)}\|_{2}+\frac{c_{k}R_{\Theta}}{m}+R_{\Theta}\|w_{k}^{(i)}\|_{2}.$ (17)
@@
$\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle\\ \leqslant\frac{R_{\Theta}}{m}\sum_{k=1}^{t}\eta_{k}c_{k}+MR_{\Theta}\sum_{k=1}^{t}\eta_{k}\|v_{k}^{(i)}\|_{2}+R_{\Theta}\sum_{k=1}^{t}\eta_{k}\|w_{k}^{(i)}\|_{2}.$ (18)
@@
$\leqslant-\eta_{1}\langle L(\theta_{1}),\theta_{2}-\theta_{1}\rangle-D_{\psi}(\theta_{2},\theta_{0})+D_{\psi}(\theta^{*},\theta_{0})$
@@
$\leqslant\frac{\eta_{1}^{2}M^{2}}{2\mu}+D_{\psi}(\theta^{*},\theta_{0})+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle.$
@@
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{H_{t}}+\frac{M^{2}}{2\mu}\cdot\frac{\sum_{k=1}^{t}\eta_{k}^{2}}{H_{t}}\\ +\frac{R_{\Theta}C(t)}{m}\cdot\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
@@
$\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}+\frac{M^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}\\ +\frac{R_{\Theta}C(T)}{m}\cdot\frac{1}{\sqrt{T}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$

### v2

@@ V Corruption Reduction by Exploiting the Best of Both Worlds
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{\alpha}{2}\Big{(}1-\frac{\alpha}{M}\Big{)}^{t_{0}-1}\frac{t_{0}(t_{0}+1)}{t(t+1)}\|\theta^{*}-\theta_{0}\|_{2}^{2}\\ +R_{\Theta}t^{r}\cdot A(t,t_{0})+\frac{K^{2}}{2\alpha}+(M+1)\sqrt{p}\sigma R_{\Theta},$
@@ VI-A Least Squares Regression on Synthetic Datasets
The performance metric is the suboptimality gap $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$. The results reported are averaged over 100 independent trials. We also report the standard deviations across these 100 experiments. The bound of the total corruption until time $t$ is set to $C(t)=20t^{0.4}$. As it is difficult to find the worst case corruption pattern, we use the following procedure as a proxy to generate the corruption $\varepsilon_{i,t}$ at each worker $i\in[m]$ and at each time $t$. In particular, let $\Gamma(\tau)=\sum_{k=1}^{\tau}c(k)^{2}$ denote the corruption budget used up to time and including $n$. Then, the available corruption budget at time $t$ is $\sqrt{C(t)^{2}-\Gamma(t-1)}$. Then, we distribute this available corruption budget to each worker randomly. The sign of adversarial corruption at each worker $\operatorname{sign}(\varepsilon_{i,t})$ is set to $-\operatorname{sign}(g^{\prime}_{i,t})$ for all $i\in[m]$. Also, the stepsize is set to $\eta_{t}=1/\sqrt{t}$.
@@
Since the effectiveness of the mirror descent procedure relies on the choice of the mirror map $\psi$, we conducted additional experiments to evaluate the performance of mirror descent under $\psi$’s. Specifically, we considered the negative entropy and the $\ell_{1}$-norm as mirror maps, in addition to the $\ell_{2}$-norm squared used in the preceding experiments. The results, presented in Fig. 4, demonstrate that the quadratic mirror map is the most suitable for least squares regression.
Figure 4: Performances of different mirror maps $\psi$ for least squares regression
@@
As shown in Fig. 5, if one chooses the stepsize sequence $\eta_{k}=\frac{\alpha}{M}H_{k}$, the adversarial corruption accumulates after a certain point, resulting in a stagnation of the suboptimality gap. However, it is also worth noting that the gap decreases much more rapidly when $\eta_{k}=\frac{\alpha}{M}H_{k}$ compared to when $\eta_{k}=\frac{2}{k+1}H_{k}$ initially. In our proposed algorithm with a restart mechanism RDGD-Restart, the gap decreases rapidly until a transition time $t_{0}$ which can be determined analytically. After time $t_{0}+1$, we switch the stepsize sequence from $\eta_{k}=\frac{\alpha}{M}H_{k}$ to $\eta_{k}=\frac{2}{k+1}H_{k}$. As can be seen, the gap subsequently decreases with $t$. Our proposed algorithm RDGD-Restart converges faster than the $\eta_{k}=\frac{2}{k+1}H_{k}$ case throughout the entire process, indicating a smaller time complexity to achieve the same suboptimality gap. The experimental results presented in Fig. 5 thus corroborate our theoretical analysis in Section V. Hence, our proposed algorithm with a restart mechanism RDGD-Restart  successfully takes advantage of the exponential decrease in the gap until corruption accumulation occurs, then mitigates the adverse effects of the accumulation by switching to another stepsize sequence.
@@ VI-B Classification using L2 Support Vector Machines
As shown in Fig. 6, RDGD converges rapidly and achieves high accuracy at any level of the corruption budget. However, the accuracy of the classification of the vanilla DGD algorithm is lower than RDGD and is highly sensitive to the level of the corruption budget. DGD fails when the corruption budget exceeds a certain level (e.g., $r=0.35$). In Fig. 7, we display the decision boundaries for the first two dimensions of the parameter vectors $\hat{\theta}_{\rm{RDGD}}$ and $\hat{\theta}_{\rm{DGD}}$ obtained by RDGD and DGD respectively when $r=0.35$. Note that $\hat{\theta}_{\rm{RDGD}}$ is almost identical to $\theta^{*}$. However, $\hat{\theta}_{\rm{DGD}}$ deviates significantly from $\theta^{*}$, showing that DGD is not robust.
@@ VI-C Softmax Classification on the MNIST Dataset
VI-C1 Performance under a bounded amount of adversarial corruptions
@@
In Fig. 8, we compare RDGD with vanilla DGD under varying levels of corruption budgets $C(T)$, with a fixed noise variance $\sigma^{2}=0.5$ in the channels. Specifically, the total corruption bound up to time $T$ is set as $C(T)=100T^{r}$, where $r\in\{0.2,0.3,0.4\}$. The total corruption budget is also uniformly allocated to each iteration and each worker as in the previous subsection. As shown in Fig. 8, RDGD  outperforms the traditional DGD algorithm across all levels of the corruption budget. Furthermore, we observe that RDGD achieves higher accuracy as the value of $r$ and hence, the amount of corruption, decreases. The vanilla DGD algorithm reduces to random guessing (success rate $\approx 10\%$) when the corruption budget exceeds a certain level (i.e., $r=0.4$).
To further evaluate the performance of RDGD, we conducted experiments comparing it against three baseline algorithms: the traditional DGD algorithm, Trimmed Mean by Yin et al. [12], and Krum by Blanchard et al. [14]. The total corruption bound up to time $T$ and the fraction of Byzantine workers are set to be $C(T)=150T^{0.3}$ and $0.3$ respectively. Note that the baseline algorithms (Trimmed Mean and Krum) are only applicable when the fraction of Byzantine workers is less than half, whereas RDGD can handle an arbitrary proportion. The total corruption budget is uniformly allocated to each iteration and each Byzantine worker. The results, shown in Fig. 9, indicate that RDGD consistently outperformed the baseline methods over all iterations. RDGD achieved a faster convergence rate and accuracy compared to DGD, Trimmed Mean, and Krum.
Figure 9: Accuracy rates of different algorithms for the MNIST dataset with adversarial corruption
VI-C2 Performance under adversarial corruption in the final iterations
As mentioned, it is difficult to find the worst case corruption pattern. In addition to allocating the available corruption budget to each iteration and each Byzantine worker uniformly, we also evaluate the performance of RDGD and the baselines when the corruption is only allocated to the last $20\%$ of iterations; this is particularly challenging for all algorithms as they have to suddenly adapt to the presence of the corruptions towards the end of the horizon. The total corruption bound up to time $T$ and the fraction of Byzantine workers are set to $C(T)=150T^{0.3}$ and $0.3$ respectively. Fig. 10 illustrates the performance of different algorithms when the corruption is introduced only during the last 20% of the iterations (i.e., from $161$ to $200$). In the first $160$ iterations, when no corruption is present, all algorithms exhibit steady increases in accuracy, with RDGD demonstrating the fastest convergence and the highest accuracy. However, in the last 20% of the iterations, the malicious gradients degrade the performance of the baselines, causing noticeable drops in their accuracy, whereas RDGD maintains its robustness.
Figure 10: Accuracy rates for different algorithms for the MNIST dataset with allocating corruptions to last $20\%$ of iterations
VI-C3 Performance under periodic adversarial corruption
To further show it is difficult to find the worst-case corruption pattern, we consider periodic adversarial corruptions, where the corruption is introduced at the $[50,100,150,200]$-th iterations. The total corruption bound up to time $T$ and the fraction of Byzantine workers are set to $C(T)=150T^{0.3}$ and $0.3$ respectively. Hence, the corruption level at these $4$ iterations is large given that we keep the corruption level the same at $C(T)=150T^{0.3}$. The sign of adversarial corruption at each Byzantine worker $\operatorname{sign}(\varepsilon_{i,t})$ is set to be $-\operatorname{sign}(\tilde{g}_{i,t})$. Then we add the corruption vector $\varepsilon_{i,t}$ to $g^{\prime}_{i,t}$, and obtain $\bar{g}_{i,t}=g^{\prime}_{i,t}+\varepsilon_{i,t}$. We checked the signs of each of the dimensions of $\bar{g}_{t}\in\mathbb{R}^{p}$ and notice that for each Byzantine worker, more than $90\%$ of the dimensions of $\bar{g}_{t}$ have their signs flipped. This shows that the corruption added significantly alters the direction of the gradient vector $\bar{g}_{t}$, implying that its effect is substantial. As shown in Fig. 11, DGD exhibits a distinctive “sawtooth” pattern, sharp periodic drops in accuracy, followed by gradual recovery. These periodic drops occur when corruptions are introduced, demonstrating DGD’s vulnerability to “sudden” Byzantine attacks. Meanwhile, the performance of RDGD remains high throughout the whole process. This experiment again demonstrates RDGD’s robustness.
Figure 11: Accuracy rates for different algorithms for the MNIST dataset with periodic corruption
VI-C4 Performance under “A Little Is Enough” attack
Besides adversarial corruption with a fixed budget, we also evaluate RDGD for the “A Little Is Enough” (ALIE) attack proposed by Baruch et al. [34], which is designed to destroy robust aggregation methods such as Trimmed Mean and Krum. We denote the set of Byzantine workers as $\mathcal{B}$. ALIE attacks the gradients at the Byzantine worker $i\in\mathcal{B}$ by setting their updates to be $g_{i,t}=\mu_{t}+z\cdot\sigma_{t}$, where $\mu_{t}$ and $\sigma_{t}$ are respectively the mean and standard deviation of the gradients at iteration $t$ among all the $m$ workers, and $z$ is an attack intensity parameter. This allows adversarial updates to remain statistically plausible, making them extremely difficult to detect. We set $z=1.5$ and the fraction of Byzantine workers to be $0.3$.
Figure 12: Accuracy rates for different algorithms for the MNIST dataset with ALIE attack
As shown in Fig. 12, RDGD exhibited remarkable robustness against this attack, achieving a consistently accuracies. In contrast, baseline methods such as Trimmed Mean and Krum showed significant drops in accuracy after the early iterations.
@@ VII Conclusion
We may consider the following avenues for future research. First, to model the scenario in which some, but not all, workers are compromised, we may redefine $c_{t}$ in (4) as $c_{t}=\sum_{i=1}^{m}\|\varepsilon_{i,t}\|_{2}$. This encourages sparsity among the $\varepsilon_{i,t}$’s for fixed $t$. Second, we may derive information-theoretic lower bounds on the suboptimality gap as a function of the corruption $C(T)$ to determine the tightness of the upper bounds.
@@ Appendix B Proof of Lemma 2
$H_{t}L_{t}-H_{t-1}L_{t-1}$ $\geqslant\eta_{t}L(\theta_{t})-\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t}\rangle$
@@
$=\eta_{t}L(\theta_{t})-\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle$
$\;+\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+D_{\psi}(\theta_{t+1},\theta_{t}).$
@@
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})-\eta_{t}L(\theta_{t})\\ +\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\\ -D_{\psi}(\theta_{t+1},\theta_{t}).$
From RDGD in Algorithm 1 and the convexity of $L$, we can write $H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})-\eta_{t}L(\theta_{t})$ as:
@@
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t})$
@@ Appendix C Proof of Theorem 1
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t}).$
By the definition of the Bregman divergence and the $\mu$-strong convexity of the mirror map $\psi$, we have $D_{\psi}(\theta_{t+1},\theta_{t})\geqslant\frac{\mu}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$. Hence,
$-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t})\\ \leqslant\eta_{t}K\|\theta_{t+1}-\theta_{t}\|_{2}-\frac{\mu}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\leqslant\frac{\eta_{t}^{2}K^{2}}{2\mu},$
where the first inequality follows from the Cauchy–Schwarz inequality and the fact that a differentiable $K$-Lipschitz continuous and convex function $L(\cdot)$ has a gradient vector $\nabla L(\cdot)$ with norm bounded by $K$ [35, Lemma 2.6], and the second inequality follows from the fact that $2ab-b^{2}\leqslant a^{2}$.
@@
$\!=\!\bigg{\langle}\frac{1}{m}\Big{(}\!\sum_{i=1}^{m}g^{\prime}_{i,k}\!+\!\sum_{i=1}^{m}\varepsilon_{i,k}\!+\!\sum_{i=1}^{m}w_{k}^{(i)}\Big{)}-\frac{1}{m}\!\sum_{i=1}^{m}g_{i,k},\theta^{*}\!-\!\theta_{k+1}\bigg{\rangle}$
$=\frac{1}{m}\bigg{\langle}\!\sum_{i=1}^{m}\!g^{\prime}_{i,k}\!-\!\sum_{i=1}^{m}\!g_{i,k},\theta^{*}\!-\!\theta_{k+1}\!\bigg{\rangle}\!+\!\frac{1}{m}\bigg{\langle}\!\sum_{i=1}^{m}\!\varepsilon_{i,k},\theta^{*}\!-\!\theta_{k+1}\!\bigg{\rangle}$
$\hskip 56.9055pt+\frac{1}{m}\bigg{\langle}\sum_{i=1}^{m}w_{k}^{(i)},\theta^{*}-\theta_{k+1}\bigg{\rangle}.$
@@
$\leqslant\frac{MR_{\Theta}}{m}\sum_{i=1}^{m}\|v_{k}^{(i)}\|_{2}+\frac{c_{k}R_{\Theta}}{m}+\frac{R_{\Theta}}{m}\sum_{i=1}^{m}\|w_{k}^{(i)}\|_{2}.$ (19)
@@
$\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle\\ \leqslant\frac{R_{\Theta}}{m}\sum_{k=1}^{t}\eta_{k}c_{k}+\frac{MR_{\Theta}}{m}\sum_{k=1}^{t}\eta_{k}\sum_{i=1}^{m}\|v_{k}^{(i)}\|_{2}\\ +\frac{R_{\Theta}}{m}\sum_{k=1}^{t}\eta_{k}\sum_{i=1}^{m}\|w_{k}^{(i)}\|_{2}.$ (20)
@@
$\leqslant-\eta_{1}\langle\nabla L(\theta_{1}),\theta_{2}-\theta_{1}\rangle-D_{\psi}(\theta_{2},\theta_{0})+D_{\psi}(\theta^{*},\theta_{0})$
@@
$\leqslant\frac{\eta_{1}^{2}K^{2}}{2\mu}+D_{\psi}(\theta^{*},\theta_{0})+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle.$
@@
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{H_{t}}+\frac{K^{2}}{2\mu}\cdot\frac{\sum_{k=1}^{t}\eta_{k}^{2}}{H_{t}}\\ +\frac{R_{\Theta}C(t)}{m}\cdot\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
@@
$\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}+\frac{K^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}\\ +\frac{R_{\Theta}C(T)}{m}\cdot\frac{1}{\sqrt{T}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$